OpenStudy (df001):

Write the expression in terms of sin and cosine and express the result in simplest form. (Step by step with distribution please.) ( I get lost while distributing "sec θ" to 1 and "cot θ") sec θ (1 + cot θ) - csc θ (1 + tan θ)

OpenStudy (df001):

This is where I get lost after distributing \[\frac{ 1 }{ \cos0 } \] to 1

OpenStudy (misty1212):

\[\frac{1}{a}(1+a)-\frac{1}{b}(1+\frac{b}{a})\]

OpenStudy (misty1212):

you get \[\frac{1}{a}+1-\frac{1}{b}-\frac{1}{a}\]after distributing

OpenStudy (misty1212):

that leaves \[1-\frac{1}{b}\]

OpenStudy (misty1212):

or \[1-\frac{1}{\sin(x)}\] or even \[1-\csc(x)\] if you prefer

OpenStudy (df001):

what are a and b

OpenStudy (misty1212):

\[a=\cos(x), b=\sin(x)\] is is 100% algebra, easier to write with \(a,b\) instead of sine and cosine

OpenStudy (df001):

oh thanks

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]